OpenStudy (anonymous):

how can i get log2 1/4 to work on a TI-83? how can i get log2 1/4 to work on a TI-83? @Mathematics

OpenStudy (anonymous):

\[\log_{2} 1/4\]

OpenStudy (lilg132):

0.5

OpenStudy (anonymous):

how do i get it to solve on a TI-83 plus graphic calc?

OpenStudy (anonymous):

Gotta use change of base formula

OpenStudy (anonymous):

If you want to use your calculator.

OpenStudy (lilg132):

if the log is 2 like u said then all you have to do is work out the square root of 1/4 or 0.25

OpenStudy (zarkon):

\[\log_{b}(a)=\frac{\ln(a)}{\ln(b)}\]

OpenStudy (anonymous):

ok.. so -2 then?

OpenStudy (anonymous):

^^ There you go!!

OpenStudy (anonymous):

nice.. that is slick.. !!!! thanks